At a 25c and 760 mm of Hg pressure a gas occupies 300 ml volume . What will be it`s pressure at a height where temperature is 10 c and volume of gas is 640 ml

Grade:9

1 Answers

Shailendra Kumar Sharma
6 years ago
PV=nRTSo for same moles of gasesPV/T = constant So P1V1/T1= P2V2/T2P2=T2*P1*V1/(T1*V2) =283*(760)*(300)/{298*640) So the pressure will be 338.4K